# A Wearable Strain-Sensor-Based Shoulder Patch for Fatigue Detection in Bicep Curls

### v1

@@ Abstract
A common challenge in home-based rehabilitation is muscle compensation induced by pain or fatigue, where patients with weakened primary muscles recruit secondary muscle groups to assist their movement, causing issues such as delayed rehabilitation progress or risk of further injury. In a home-based setting, the subtle compensatory actions may not be perceived since physiotherapists cannot directly observe patients. To address this problem, this study develops a novel wearable strain-sensor-based shoulder patch to detect fatigue-induced muscle compensation during bicep curl exercises. Built on an observation that the amplitude of a strain sensor’s resistance is correlated to the motion of a joint that the sensor is attached to, we develop an algorithm that can robustly detect the state when significant changes appear in the shoulder joint motion, which indicates fatigue-induced muscle compensation in bicep curls. The developed shoulder patch is tested on 13 subjects who perform bicep curl exercises with a 5 kg dumbell until reaching fatigue. During the experiment, the performance of the shoulder patch is also benchmarked with optical tracking sensors and surface electromyography (sEMG) sensors. Results reveal that the proposed wearable sensor and detection methods effectively monitor fatigue-induced muscle compensation during bicep curl exercises in both Real-Time and Post Hoc modes. This development marks a significant step toward enhancing the effectiveness of home-based rehabilitation by providing physiotherapists with a tool to monitor and adjust treatment plans remotely.
@@ I Introduction
Home-based rehabilitation, conducted in the patient’s residence rather than a clinical setting, can reduce these burdens but introduces its own challenges. A key concern is the lack of direct supervision by physiotherapists, making it difficult to monitor patient progress accurately. With limited functionality or fatigue, patients may unknowingly compensate during exercise, a behavior where the central nervous system recruits secondary muscles or joints to complete a movement when the primary muscle is at reduced capacity. For example, Liu et al. [3] utilized complex networks theory, showing the shifts of the muscles’ functional connection as fatigue increases, exemplifying the compensation mechanism. Although compensation helps patients to accomplish the movement, it can lead to chronic issues such as pain, reduced range of motion, and diminished capacity for motor function in the affected joints [4].
@@
In [5], sEMG signals were collected from several sEMG sensors attached to the surface of the investigated muscles. These sensors measure the electrical signal generated by the muscle fibers during muscle contraction. The joint kinematics were recorded by a motion capture system, which utilized multiple infrared cameras positioned around the capture area to detect the reflective markers attached to the key points of the object to precisely track their movement. Despite high accuracy, using sEMG sensors or motion capture systems in home-based rehabilitation is impractical due to their cost, complexity, and the need for trained professionals to operate them [6]. As a result, they have mainly been used to set up the ground truth for an exercise [7, 8] or as a validation tool of a developed sensor system [9, 10]. This highlights the necessity for more user-friendly alternatives that can capture muscle force and joint movements independently. For example, sensors such as Inertial Measurement Units (IMUs), which capture the acceleration and angular velocity [11], pressure sensors, which capture the exerted force and the applied pressure [12], or marker-free cameras, which capture the visual information of movements [13], are potential substitutes for joint kinematics detection. However, they are susceptible to environmental influences such as sensor orientation and lighting conditions. For muscle force measurement, sensing options are more limited. Although mechanomyography (MMG) [14], which measures the mechanical oscillations of muscles during contractions, provides a direct assessment, it is less commonly used compared to sEMG sensors due to less technological maturity. Dynamometers [10], while precise, are typically confined to clinical settings due to their operational complexities. In addition, it only measures the torque generated during muscle contraction against the dynamometer resistance and does not directly measure muscle force. Therefore, developing an integrated device that combines these technologies into a robust, easy-to-use system for home rehabilitation is essential. Further research is needed to enhance the accuracy, usability, and environmental resilience of these sensors, making them suitable for non-professional users in home settings.
@@
Strain gauge sensors have demonstrated versatility across various applications, including monitoring breathing rates, muscle force, and joint kinematics. Jonathan et al. [10] employed wearable soft strain sensors to track muscle deformation during contractions, discovering patterns of muscle activity similar to those detected by dynamometers. Hao et al. [15] developed a strain sensor capable of capturing joint movements, such as the bending of fingers and wrists, further illustrating the sensor’s utility in dynamic motion tracking. Chenyu et al. [16] extended the application of strain sensors to develop a choker as an interface for silent speech, employing a neural network model to analyze strain sensor signals generated during silent speech attempts. Qi et al. [17] crafted an “E-Jacket” equipped with strain sensors to monitor different postures, using a CNN-LSTM model to analyze inputs from four sensors and classify user postures including standing, walking, sitting, running, and lying.
@@
These examples highlight the broad potential of strain sensors, particularly in targeted applications for muscle activity detection and joint movement monitoring. However, these works mostly rely on arrays of strain sensors to determine joint movements via activated signals from a subset of sensors and use machine learning algorithms for complex tasks. Compared to others, strain sensors offer the advantage of providing direct measurements of the sensor formation on the skin surface, which often correlates closely with the underlying joint movement. Additionally, the strain sensors are more user-friendly due to their light weight, conformability to the skin surface, and low setup required for effective use. Co-developing medical sensors employing machine learning methods typically requires extensive pre-trained data, necessitating numerous medical trials using the developed sensor, which often involve regulatory hurdles and variability in the patients’ data, thereby increasing development costs and time. Consequently, there is a significant need to create a low-cost strain sensor system for fatigue detection, simplifying the technology and reducing reliance on extensive data sets for effective application in real-world settings.
@@
To this end, this paper developed a wearable strain-sensor-based shoulder patch to capture the onset of fatigue during bicep curl exercises. Building on the stretchable strain sensors developed by our team previously [18], we designed a shoulder patch that can be conveniently attached to a user for muscle activity monitoring. The patch is attached to the Acromioclavicular (AC) joint through kinesiology tape, and the shoulder movement is monitored by reading the strain sensor resistance. The obtained signal is processed in both Real-Time and Post Hoc modes. The former enables practical application in detecting patient’s state while exercising. The latter allows for the retrospective analysis of patients’ rehabilitation states to provide personalized rehabilitation feedback to the patient and understand their fatigue patterns.
@@
Figure 1: Overview of the study, employing the strain sensor for monitoring fatigue during bicep curl exercise. (A) Illustration of the placement of the strain sensor on the subject during the exercise. The sensor is selected to be placed on the AC joint. The rationale of the sensor placement is discussed in Sec. III-C2. (B) The working principle of the sensor developed, designed to capture deformations caused by excessive shoulder movements such as flexing and elevating, which occur as compensatory patterns during fatigue in bicep curl exercise. (C) Display of the exemplar data from the subject performing weight-free bicep curls, standard bicep curls, and fatigued bicep curls. The derivation of the data, $R_{norm}$ is discussed in Eq. (2)
@@ II-A Hardware Development
The strain sensor is fabricated using carbon nanofibers (CNFs) with Pyrograf-III, grade PR-24-XT-HHT reinforced platinum (Pt) thin films, as reported by our team’s previous work [18]. When subjected to strain, the sensor’s resistance changes with a high sensitivity and linearity. This work further adapts this sensor architecture to develop a soft wearable patch.
@@ II-A1 Fabrication of Strain Sensor
Briefly, CNFs are dispersed in isopropyl alcohol (IPA, Chem-Supply, 1 mg ml-1) with 2 wtPolyvinylpyrrolidone (PVP, Sigma, Mw $\sim$ 40,000) as a dispersing agent. The mixture is probe-sonicated for one hour to prepare the conductive ink. Simultaneously, porous polydimethylsiloxane (PDMS, Sylgard 184) substrates are prepared using a $CO_{2}$ laser cutter (Speedy 400). The ink mixture is then spray-coated on the porous PDMS thin substrate using a spray gun (Blackridge airbrush kit, needle diameter 0.2 mm, pressure 15-20 PSI) from a distance of 10 cm and then dried in an air oven at $60$\mathrm{\SIUnitSymbolCelsius}$$C for an hour.
Next, a thin layer of Pt coating is sputter-coated onto the surface of the porous CNF/PDMS substrates. Finally, the Pt/CNFs/PDMS is cut into 1 cm width and 3 cm long strips with conductive threads attached to both ends using silver paste (OK-SPI, SPI Supplies) and encapsulated with aluminum foil tape. A comprehensive illustration of the sensor manufacturing process is provided in Fig. 2.
Figure 2: The manufacturing process of the strain sensor. (1) Laser cutting the pores on the PDMS substrates. (2) Spray coating the CNF ink to the PDMS substrates. (3) Sputter coating the Pt on the CNF/PDMS substrates.
@@ II-B Software Development
Prior to the software development, preliminary experiments, as will be described in Sec. III-C3, were conducted to determine the sensor’s suitability to detect fatigue and to determine the key changes in the sensor signal associated with fatigue. These experiments concluded that placing the sensor on the AC joint is optimal for capturing fatigue-induced compensation during bicep curl exercises. Furthermore, the strain sensor effectively reflects the patient’s fatigue condition during exercise. We observed that as fatigue increases, the strain signal exhibits a visible increment in the cycle amplitude and greater signal variability. Based on this observation, we propose two algorithms: 1) Real-Time fatigue detection and 2) Post Hoc fatigue detection. The former allows the detection of whether the patient is experiencing fatigue while exercising, while the latter enables the retrospective analysis of the patient’s rehabilitation states.
@@ II-B2 Real-Time Detection
In the Real-Time mode, we aim to develop an algorithm that can identify when a subject enters the fatigue stage along with the progress of the exercise. To this end, we propose a sliding window approach, with the algorithm outlined in Algorithm 1 and the steps illustrated in Fig. 5.
First, the preprocessed sensor signals $R_{norm}$ are collected in batches of 50 data points (step ①). Each batch is then detrended (step ②; we used MATLAB’s “detrend” function in our implementation). Next, the peaks and troughs are identified in the detrended data, and the cycle amplitude is calculated (step ③). As each batch is processed, the minimum cycle amplitude is recorded and used as a reference for comparison (step ④). A fatigue threshold, $\tau$ of 3.5 is established by considering the linear relationship between the shoulder kinematics and the strain sensor that will be discussed in Sec. III-C2 and the increment of the shoulder elevation ROM between standard bicep curls and fatigued bicep curls [5]. For each batch, the ratio of the current cycle amplitude to the reference cycle amplitude is calculated (step ⑤). If this ratio exceeds the predefined threshold for two or more consecutive batches, the start time of the first batch in the sequence is recorded as the fatigue time (step ⑥), denoted as $t_{r}$. If the condition in step ⑥ is not satisfied before the subject completes the exercise, $t_{r}$ is set to be the end time of the signal, indicating the fatigue is not detected as the fatigue-induced compensation may not occur.
@@
  Output: Fatigue Timestamp $t_{r}$, Fatigue Flag $fat$
@@ II-B3 Post Hoc Detection
  Output: Fatigue Timestamp $t_{p}$, Fatigue Flag $fat$
@@
  $t_{1},fat\leftarrow$ Compute Real Time Fatigue Detection :RealTime($R_{raw}$,$\overline{R}_{static}$,$n$,$\tau$) {step ①}
@@
  $filteredPeaks\leftarrow$ Apply time-based filter to combine or remove noises in $peaks$, {step ④}
  $t_{2}\leftarrow$ Output the time of the first peak in $filteredPeaks$step ⑤}
@@ III Experiment and Validation
In this section, we first detail the experimental setup and discuss the protocols employed in this study, including the bicep curl and fatigue protocols. Next, we outline the experiments conducted to determine the suitability of the sensors for detecting fatigue and to identify changes in the sensor signal associated with fatigue. This is followed by a detailed description of the subject study. Finally, we relate the findings from our previous study on fatigue during bicep curl to the proposed strain sensor and outline methods used to process data from different sensor systems to validate the sensor’s efficiency.
@@ III-A Experiment Setup
Figure 7: Two joint angles during bicep curl in this study. The red dotted lines represent the references; the purple line links the clavicle and the acromioclavicular joint; the blue and green lines depict the upper arm limb, and the lower arm limb respectively. The arrows indicate the positive directions.
@@ III-B2 Subject’s Fatigue Protocol
To assess the subject’s perceived exertion and fatigue during the study, we employ the Borg CR-10 Scale [21]. During the experiment, subjects evaluate themselves based on the Rated Perceived Exertion (PRE), with a level of seven as the threshold of fatigue. Fatigue is defined either by the inability to complete a bicep curl cycle or verbal confirmation of the inability to continue. At that point, the subject’s perceived onset of fatigue, marked by timestamp, $t_{s}$, is recorded.
@@ III-C Preliminary Experiments
Three preliminary experiments were conducted to evaluate the suitability of the strain sensor for detecting changes in kinematics, sEMG signals, and fatigue. It is important to note that the sensor placements in these preliminary experiments were different from those described in Section III-A, which was specific to the subject study.
@@
III-C1 Experiment 1: Comparison of Changes in Strain Resistance Against sEMG

### v2

@@ Abstract
A common challenge in home-based rehabilitation is muscle compensation induced by pain or fatigue, where patients with weakened primary muscles recruit secondary muscle groups to assist their movement, causing issues such as delayed rehabilitation progress or risk of further injury. In a home-based setting, the subtle compensatory actions may not be perceived since physiotherapists cannot directly observe patients. To address this problem, this study develops a novel wearable strain-sensor-based shoulder patch to detect fatigue-induced muscle compensation during bicep curl exercises. Built on an observation that the amplitude of a strain sensor’s resistance is correlated to the motion of a joint that the sensor is attached to, we develop an algorithm that can robustly detect the state when significant changes appear in the shoulder joint motion, which indicates fatigue-induced muscle compensation in bicep curls. The developed shoulder patch is tested on 13 subjects who perform bicep curl exercises with a 5 kg dumbbell until reaching fatigue. During the experiment, the performance of the shoulder patch is also benchmarked with optical tracking sensors and surface electromyography (sEMG) sensors. Results reveal that the proposed wearable sensor and detection methods effectively monitor fatigue-induced muscle compensation during bicep curl exercises in both Real-Time and Post Hoc modes. This development marks a significant step toward enhancing the effectiveness of home-based rehabilitation by providing physiotherapists with a tool to monitor and adjust treatment plans remotely.
@@ I Introduction
Home-based rehabilitation, conducted in the patient’s residence rather than a clinical setting, can reduce these burdens but introduces its own challenges. A key concern is the lack of direct supervision by physiotherapists, making it difficult to monitor patient progress accurately. A crucial part of effectively monitoring patient progress in home-based rehabilitation is the ability to gauge patient’s effort. Substantial effort can quickly lead to fatigue, and assessing it has been challenging. In clinical rehabilitation, physiotherapists employ various methods, including subjective measures such as Borg Rate of Perceived Exertion (RPE) [3], and objective measurements via physiological signal monitoring such as heart rate, EMG signals, and performance-based tests such as range of motion, force, speed, and completion of a task. However, fatigue is multifaceted, complex, and manifests differently across individuals. As a result, assessing fatigue is challenging and requires instinct from the healthcare provider.
With fatigue, patients may unknowingly compensate during exercise, a behavior where the secondary muscles or joints are increasingly engaged to complete a movement when the primary muscle is at reduced capacity. For example, Liu et al. [4] utilized complex networks theory, showing the shifts of the muscles’ functional connection as fatigue increases, exemplifying the compensation mechanism. Although compensation helps patients accomplish the movement, it can lead to chronic issues such as pain, reduced range of motion, and diminished capacity for motor function in the affected joints [5].
@@
In [6], sEMG signals were collected from several sEMG sensors attached to the surface of the investigated muscles. These sensors measure the electrical signal generated by the muscle fibers during muscle contraction. The joint kinematics were recorded by a motion capture system, which utilized multiple infrared cameras positioned around the capture area to detect the reflective markers attached to the key points of the object to precisely track their movement. Despite high accuracy, using sEMG sensors or motion capture systems in home-based rehabilitation is impractical due to their cost, complexity, and the need for trained professionals to operate them [7]. As a result, they have mainly been used to set up the ground truth for an exercise [8, 9] or as a validation tool of a developed sensor system [10, 11]. This highlights the necessity for more user-friendly alternatives that can capture muscle force and joint movements independently. For example, sensors such as Inertial Measurement Units (IMUs), which capture the acceleration and angular velocity [12, 13], pressure sensors, which capture the exerted force and the applied pressure [14], or marker-free cameras, which capture the visual information of movements [15, 16], have been maturely used for joint movement detection, particularly in rehabilitation settings.
However, they are susceptible to environmental influences such as sensor orientation and lighting conditions. For muscle force measurement, sensing options are more limited. Although mechanomyography (MMG) [17], which measures the mechanical oscillations of muscles during contractions, provides a direct assessment, it is less commonly used compared to sEMG sensors due to less technological maturity. Dynamometers [11], while precise, are typically confined to clinical settings due to their operational complexities. In addition, it only measures the torque generated during muscle contraction against the dynamometer resistance and does not directly measure muscle force. Therefore, developing an integrated device that combines these technologies into a robust, easy-to-use system for home rehabilitation is essential. Further research is needed to enhance the accuracy, usability, and environmental resilience of these sensors, making them suitable for non-professional users in home settings.
@@
Strain gauge sensors have demonstrated versatility across various applications, including monitoring breathing rates, muscle force, and joint kinematics. Jonathan et al. [11] employed wearable soft strain sensors to track muscle deformation during contractions, discovering patterns of muscle activity similar to those detected by dynamometers. Sun et al. [18] developed a strain sensor capable of capturing joint movements, such as the bending of fingers and wrists, further illustrating the sensor’s utility in dynamic motion tracking. Tang et al. [19] extended the application of strain sensors to develop a choker as an interface for silent speech, employing a neural network model to analyze strain sensor signals generated during silent speech attempts. Lin et al. [20] crafted an “E-Jacket” equipped with strain sensors to monitor different postures, using a CNN-LSTM model to analyze inputs from four sensors and classify user postures including standing, walking, sitting, running, and lying.
@@
These examples highlight the broad potential of strain sensors, particularly in targeted applications for muscle activity detection and joint movement monitoring. However, these works mostly rely on arrays of strain sensors to determine joint movements via activated signals from a subset of sensors and use machine learning algorithms for complex tasks. Compared to others, strain sensors offer the advantage of providing direct measurements of the sensor formation on the skin surface, which often correlates closely with the underlying joint movement. Additionally, the strain sensors are more user-friendly due to their lightweight, conformability to the skin surface, and low setup required for effective use.
Co-developing medical sensors employing machine learning methods typically requires extensive pre-trained data, necessitating numerous medical trials using the developed sensor, which often involve regulatory hurdles and variability in the patients’ data, thereby increasing development costs and time. Consequently, there is a significant need to create a low-cost strain sensor system for fatigue detection, simplifying the technology and reducing reliance on extensive data sets for effective application in real-world settings.
@@
To this end, this paper introduces a wearable strain-sensor-based shoulder patch to capture the onset of fatigue during bicep curl exercises. Building on the stretchable strain sensors developed by our team previously [21], we designed a shoulder patch that can be conveniently attached to a user for muscle activity monitoring. The patch is attached to the Acromioclavicular (AC) joint through kinesiology tape, and the shoulder movement is monitored by reading the strain sensor resistance. The obtained signal is processed in both Real-Time and Post Hoc modes. The former enables practical application in detecting a patient’s state while exercising. The later allows for the retrospective analysis of patients’ rehabilitation states to provide personalized rehabilitation feedback to the patient and understand their fatigue patterns.
@@
Figure 1: Overview of the study, employing the strain sensor for monitoring fatigue during bicep curl exercise. (A) Illustration of the placement of the strain sensor on the subject during the exercise. The sensor is selected to be placed on the AC joint. The rationale of the sensor placement is discussed in Sec. III-C3. (B) The working principle of the sensor developed, designed to capture deformations caused by excessive shoulder movements such as flexing and elevating, which occur as compensatory patterns during fatigue in bicep curl exercise. (C) Display of the exemplar data from the subject performing weight-free bicep curls, standard bicep curls, and fatigued bicep curls. The derivation of the data, $R_{norm}$ is discussed in Eq. (2)
@@ II-A Hardware Development
The strain sensor is fabricated using carbon nanofibers (CNFs) with Pyrograf-III, grade PR-24-XT-HHT reinforced platinum (Pt) thin films, as reported by our team’s previous work [21]. Specifically, the sensor demonstrates a gauge factor of approximately 417, a linear response rate of 97%, a wide sensing range up to 121% of strain, and excellent durability, maintaining stable performance over 2000 cycles of stretching and releasing. These features make it well-suited for detecting fatigue-induced compensatory movements. This work further adapts this sensor architecture to develop a soft wearable patch.
@@ II-A1 Fabrication of Strain Sensor
[…]
Figure 2: The manufacturing process of the strain sensor. (1) Fabricating the PDMS substrates. (2) Spray coating the CNF ink onto the PDMS substrates. (3) Sputter coating the Pt on the CNF/PDMS substrates.
@@ II-A3 Sensor Interface Setup
II-A4 Cost Analysis
The primary components of the sensor including the CNFs, Pt, and PDMS, incur a material cost of approximately \$\\$0.05 USD per unit. Additional materials that are needed to assemble the sensor such as the silver paste, aluminum foil, kinesiology tape, and conductive wire, raise the total to approximately \$\\$0.74 USD per unit. All the materials may require replacement if they degrade over time. Meanwhile, the electronics hardware, such as the Arduino UNO, parrot hook, resistor, and breadboard, which is generally not subject to frequent replacement, adds a deployment cost of about \$\\$34.77 USD per setup.
@@ II-B Software Development
Prior to the software development, preliminary experiments, as will be described in Sec. III-C4, were conducted to determine the sensor’s suitability to detect fatigue and to determine the key changes in the sensor signal associated with fatigue. These experiments concluded that placing the sensor on the AC joint is optimal for capturing fatigue-induced compensation during bicep curl exercises. Furthermore, the strain sensor effectively reflects the patient’s fatigue condition during exercise. We observed that as fatigue increases, the strain signal exhibits a visible increment in the cycle amplitude and greater signal variability. Based on this observation, we propose two algorithms: 1) Real-Time fatigue detection and 2) Post Hoc fatigue detection. The former allows the detection of whether the patient is experiencing fatigue while exercising, while the latter enables the retrospective analysis of the patient’s rehabilitation states.
@@ II-B2 Real-Time Detection
In the Real-Time mode, we aim to develop an algorithm that can identify when a subject enters the fatigue stage, along with the progress of the exercise. To this end, we propose a sliding window-based algorithm outlined in Algorithm 1 and the steps illustrated in Fig. 5.
First, the preprocessed sensor signals $R_{norm}$ are collected in batches of 50 data points for approximately 2s. This configuration is chosen due to an empirical observation that a bicep curl cycle takes approximately 2 seconds (step ①). Each batch is then detrended (step ②; we used MATLAB’s “detrend” function in our implementation). Next, the peaks and troughs are identified in the detrended data, and the cycle amplitude is calculated (step ③). As each batch is processed, the minimum cycle amplitude is recorded and used as a reference for comparison (step ④). A fatigue threshold $\tau$ of 3.5 is established by considering the linear relationship between the shoulder kinematics and the strain sensor that will be discussed in Sec. III-C3 and the increment of the shoulder elevation ROM between standard bicep curls and fatigued bicep curls [6]. For each batch, the ratio of the current cycle amplitude to the reference cycle amplitude is calculated (step ⑤). If this ratio exceeds the predefined threshold for two or more consecutive batches, the start time of the first batch in the sequence is recorded as the fatigue time (step ⑥), denoted as $t_{r}$. If the condition in step ⑥ is not satisfied before the subject completes the exercise, $t_{r}$ is set to be the end time of the signal, indicating the fatigue is not detected as the fatigue-induced compensation may not occur.
@@
Output: Fatigue Timestamp $t_{r}$, Fatigue Flag fatfat
@@ II-B3 Post Hoc Detection
Output: Fatigue Timestamp $t_{p}$, Fatigue Flag fatfat
@@
[…]
@@
$filteredPeaks\leftarrow$ Apply time-based filter to combine or remove noises in peakspeaks, {step ④}
$t_{2}\leftarrow$ Output the time of the first peak in filteredPeaksfilteredPeaks {step ⑤}
@@ III Experiment and Validation
In this section, we first describe the experimental setup and discuss the protocols employed in this study, including the bicep curl and fatigue protocols. Next, we outline the experiments conducted to determine the suitability of the sensors for detecting fatigue and to identify changes in the sensor signal associated with fatigue. This is followed by a detailed description of the subject study. Finally, we relate the findings from our previous study on fatigue during bicep curl to the proposed strain sensor and outline methods used to process data from different sensor systems to validate the sensor’s efficiency.
@@ III-A Experiment Setup
Figure 7: Two joint angles during bicep curl in this study. The red dotted lines represent the references; the purple line links the clavicle and the acromioclavicular joint; the blue and green lines depict the upper arm limb and the lower arm limb, respectively. The arrows indicate the positive directions.
@@ III-B2 Subject’s Fatigue Protocol
To assess the subject’s perceived exertion and fatigue during the study, we employ the Borg CR-10 Scale [24]. During the experiment, subjects self-evaluate based on their RPE, with a level of seven as the threshold of fatigue. Fatigue is defined either by the inability to complete a bicep curl cycle or verbal confirmation of the inability to continue. At that point, the subject’s perceived onset of fatigue, marked by timestamp, $t_{s}$, is recorded.
@@ III-C Preliminary Experiments
Four preliminary experiments were conducted to evaluate the suitability of the strain sensor for detecting changes in kinematics, sEMG signals, and fatigue. It is important to note that the sensor placements in these preliminary experiments were different from those described in Section III-A, which was specific to the subject study.
@@
III-C1 Experiment 1: Static-Noise and Power Consumption Assessment of Strain Sensor
This experiment aimed to quantify the sensor’s baseline static noises and additional noise when the system is disturbed, recognizing the signal acquired from human skin, a highly deformable surface and the signal quality is constrained by the acquisition system. The experimental setup used was identical to that illustrated in Fig. 8, excluding the sEMG sensor and Motive markers. The outputs from the strain sensor were recorded while the subject remained static under two conditions, where the interconnecting cables remained undisturbed and when the interconnecting cables were deliberately jostled to emulate the worst-case handling. As illustrated in Fig. 10, our application reported a signal-to-noise ratio ranging from 18.0 dB to 28.1 dB, indicating a consistently clean signal. The result also indicated no reduction in the signal-to-noise ratio when the cable is vigorously manipulated. This preliminary study concluded that as long as the parrot hook is secured on the subject’s shoulder, it does not affect the functionality of the strain sensor.
Figure 10: Comparison of the Normalized Strain Resistance under two test conditions. The blue graph represents the condition where the cable is left undisturbed. The red graph represents the condition where the cable is vigorously oscillated. The signal-to-noise ratio remains similar, indicating that cable motion does not degrade the signal quality.
This experiment also quantifies the nominal power consumption of the circuit. The strain sensor forms a voltage divider circuit with a 1k $\Omega$ reference resistor, $(R)$ and it is excited from a regulated 5.0 V from Arduino, $(V_{in})$. At static, the nominal strain resistance, $(R_{strain})$ is recorded at approximately 700 $\Omega$. Using Equation 3, the circuit is dissipating 14.7 mW.
$P=\frac{V_{in}^{2}}{R+R_{strain}}$ (3)
III-C2 Experiment 2: Comparison of Changes in Strain Resistance Against sEMG